OpenStudy (anonymous):

elastic ball (m=0.1kg) fall from 2m touch the rigid floor and bounce to 1.6m. how to calculate the momentum of the ball before and after the choc?

OpenStudy (anonymous):

Momentum is given by the product of a body's mass and velocity. Mass is given to be 0.1 kg, so we calculate velocity just before and just after impact using the formula \[v^{2}=u ^{2}+2as\] where v is final velocity u is initial velocity a is acceleration acting on body (here, a=g=9.8 m/s^2 since acceleration is only due to gravity) s is the distance travelled Velocity of ball just before it hits the floor: Ball started from rest, u=0 m/s, and travelled s=2m Plugging all of these in the formula, we get final velocity \[v=2\sqrt{10} m/s\]\[v=6.33m/s\] Momentum \[p=mv=(0.1)(6.33)=0.63kg.m/s\] Velocity of ball just after impact: Ball bounces to a height s=1.6 m before momentarily coming to rest (final velocity v=0) Here we calculate initial velocity u \[u ^{2}=v ^{2}-2as=0-(-10)(1.6)=16\]\[u=4m/s\] Momentum just after the ball hits the floor: \[p=m u=(0.1)(4)=0.4kg.m/s\]

OpenStudy (anonymous):

(**Considered g=10m/s^2, sorry about that.)